Provide the missing data in the following table for a distributor of martial arts products: (Enter Turnover and ROI answe
Answer #1

Alpha:

Sales = turnover*Average operating assets

= 5*439000 = $2195000

Net operating income = 2195000*8% = 175600

Return on investment = 175600/2195000 = 8%

Bravo:

Margin = 49200/307500 = 16%

Average operating assets = 49200/0.40 = $123000

Turnover = 307500/123000 = 2.5

Charlie:

Sales = 63900/0.15 = $426000

Average operating assets = 63900/0.30 = $213000

Turnover = 426000/213000 = 2